OpenStudy (anonymous):

Find the height of a person who is 1.7 m tall, in feet and inches, rounded to the nearest inch (NOTE: 1 foot = 12 inches). ft. in.

OpenStudy (fibonaccichick666):

so on the last problem you saw how dan multiplied by something equivalent to 1 to change the units, can you figure out what that is here?

OpenStudy (danjs):

lol sorry fibonacci, you can take this one

OpenStudy (fibonaccichick666):

can you erase your comment then? It sort of gives the answer away

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

thanks :)

OpenStudy (fibonaccichick666):

so, like I was saying, you need something called the conversion factor, this is a fraction that is equivalent to 1 which goes between the units you are working with. Do you know it for m to cm?

OpenStudy (anonymous):

yes we muliply

OpenStudy (fibonaccichick666):

by what? What is our conversion factor?

OpenStudy (anonymous):

100

OpenStudy (fibonaccichick666):

for meters to centimeters we multiply by 100?

OpenStudy (anonymous):

no we multiply by 10

OpenStudy (fibonaccichick666):

I don't think we do that either

OpenStudy (fibonaccichick666):

how many cm are in a meter?

OpenStudy (anonymous):

1 cm in 100 meter

OpenStudy (fibonaccichick666):

A cm is bigger than a meter?

OpenStudy (anonymous):

no sorry cm is smaller 100 meter in 1 cm

OpenStudy (fibonaccichick666):

you said the same thing again

OpenStudy (anonymous):

cm is smaller than meter

OpenStudy (fibonaccichick666):

that is correct, but you need the number of cm in a meter

OpenStudy (anonymous):

ok which number

OpenStudy (fibonaccichick666):

well, you should know this. if you do not, a google search will bring it up very quickly.

OpenStudy (fibonaccichick666):

what does 'centi' stand for?

OpenStudy (danjs):

EX] If you have 6 apples, and 8 oranges equals 3 apples, how many oranges do you have?

OpenStudy (fibonaccichick666):

how many 'cents' are in a dollar?

OpenStudy (danjs):

\[\frac{ 6 apples}{ 1 }\frac{ 8 oranges }{ 3 apples }\]

OpenStudy (anonymous):

100 cents

OpenStudy (fibonaccichick666):

ok so what does cent mean?

OpenStudy (anonymous):

18/ 8

OpenStudy (anonymous):

100 cents in one dollar

OpenStudy (fibonaccichick666):

so centi stands for \(\frac{1}{100}\)

OpenStudy (fibonaccichick666):

therefore, how many centimeters are in a meter?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

am i right?

OpenStudy (fibonaccichick666):

that is correct. 100cm in 1 m

OpenStudy (fibonaccichick666):

now, that is your first conversion factor.

OpenStudy (fibonaccichick666):

the next one you need is how many centimeters are in an inch

OpenStudy (anonymous):

2.54

OpenStudy (fibonaccichick666):

in a sentence please

OpenStudy (anonymous):

there is 2.54 centi in 1 inch

OpenStudy (fibonaccichick666):

ok so now, how about inches to feet?

OpenStudy (anonymous):

12 inches in one foot

OpenStudy (fibonaccichick666):

good, so now, you have all of your conversion factors it is time to set up your dimensional analysis. we start with \[ \frac{1.7m}{1} \times \frac{?}{?} \times \frac{?}{?} \times \frac{?ft}{?}\]

OpenStudy (fibonaccichick666):

so, you need to set up your conversion factors so that the units cancel out and you are left with feet. Please do so

OpenStudy (anonymous):

what goes where

OpenStudy (fibonaccichick666):

that's for you to decide.

OpenStudy (fibonaccichick666):

You need your units to cancel.

OpenStudy (anonymous):

1.7m/1 = 170m

OpenStudy (anonymous):

i got my answer as 5 feet 7 inches

OpenStudy (anonymous):

am i right?

OpenStudy (fibonaccichick666):

no, how did you get that?

OpenStudy (anonymous):

1.7 meters is 5.58 feet or 66.93 inches 12 inches = 1 foot.

OpenStudy (fibonaccichick666):

please use the fractions I have set up

OpenStudy (fibonaccichick666):

wait I'm sorry, that may be correct

OpenStudy (fibonaccichick666):

I was reading that as 170 meters for some reason

OpenStudy (fibonaccichick666):

oh that is because you said 1.7 m=170?

OpenStudy (fibonaccichick666):

but anyways, your answer is correct but I have no clue how you did it since you didn't show your work.

OpenStudy (anonymous):

so am i right? 5'7''

OpenStudy (fibonaccichick666):

I wish you would show the work, but yes